Supposed repetition are not allowed.
a) How many three-digit numbers can be formed from the six digits 2, 3, 5, 6, 7 and 9?
b) How many are less than 500? [2]
c) How many are even? [2]
3.2 Simplify
(
1
Expert's answer
2021-09-20T11:42:19-0400

a) There are six digits 2, 3, 5, 6, 7 and 9. Supposed repetition are not allowed.

Then


"P(6,3)=\dfrac{6!}{(6-3)!}=6(5)(4)=120"

120 three-digit numbers can be formed.


(b)

Digit 2 is at the first place


"P(5,2)=\dfrac{5!}{(5-2)!}=5(4)=20"

Digit 3 is at the first place


"P(5,2)=\dfrac{5!}{(5-2)!}=5(4)=20"

"20+20=40"

40 three-digit numbers less 500 can be formed.


(c) Digit 2 is at the third place


"P(5,2)=\dfrac{5!}{(5-2)!}=5(4)=20"


Digit 6 is at the third place


"P(5,2)=\dfrac{5!}{(5-2)!}=5(4)=20""20+20=40"

40 three-digit even numbers can be formed.